A solution of tartaric acid is prepared of concentration 1.2174 g L -1 and it is found that 50.0 ml of the solution requires 21.48 mL of 3.78 × 10-2 mol L -1 sodium hydroxide for complete neutralization. The formula of tartaric acid is C4H6O6 and its relative molecular mass is 150.2. Calculate how many of the hydrogens in a molecule of tartaric acid are ionised in aqueous solution.
Answer #1

Mass of acid = 50×1.2174/1000=> 60.87/1000 g

Gram equivalent of Acid= gram equivalent of base

Mole ×n=M×V(L)

60.87×n/1000×150.2 = 21.48×0.0378/1000

n = 2

So 2 hydrogens in a molecule of acid is ionized